OpenStudy (anonymous):

The sum of the roots of the equation, x^2 + | 2x - 3|– - 4 = 0, is :- (1) - 2 (2) -–(2)^1/2 (3) –(2)^1/2 (4) 2

OpenStudy (anonymous):

i think there is some problem with the question because when i sub in all the answers, non of them equal to 0

OpenStudy (anonymous):

but the solution should be like this | 2x - 3|–=4-x^2 [(2x-3)^2]^(1/2)=4-x^2 (2x-3)^2=(4-x^2)^2 expand it and it will become x^4-4x^2+12x+7=0

OpenStudy (anonymous):

Options you are substituting are sum of roots not the roots of equation ! Answer is C \[\sqrt{2}\] but dont know how :(

OpenStudy (anonymous):

should have some problem with the question because if sub 2^1/2 the answer is not equal to 0 as well or you can just mark it and ask your tutor tomorrow =]

OpenStudy (anonymous):

you can't substitute 2^1/2 because it is sum of roots not the root of equation !